OpenStudy (anonymous):

A car can travel 476 miles on 14 gallons of gas. Write an equation relating the distance d to the number of gallons g. How many gallons of gas does this car need to travel 578 miles?

OpenStudy (anonymous):

476/14=d/g 476/14=34 34 miles per gallon 578/34 is 17 17 gallons